OpenStudy (anonymous):

N/8 + 6 = 58; N equals

OpenStudy (ash2326):

We have \[\frac N 8+6=58\] Multiply both sides by 8 \[N+8\times6=58\times 8\] Can you solve now?

OpenStudy (jiteshmeghwal9):